# Geometric measures of CHSH quantum nonlocality: characterization, quantification, and comparison by distances and operations

### v4

@@ V.1 Bell-diagonal states
$\rho_{BD}=\frac{1}{4}\bigg{[}\mathcal{I}_{2}\otimes\mathcal{I}_{2}+\displaystyle\sum_{i=1}^{3}a_{i}\sigma_{i}\otimes\sigma_{i}\bigg{]}\,,$ (39)
where $\mathcal{I}_{2}$ denotes the identity matrix in two dimensions. Matrix $\alpha$ of Eq. (32) can then be recast in the form
$\alpha_{BD}=\begin{bmatrix}1&0&0&0\\ 0&a_{1}&0&0\\ 0&0&a_{2}&0\\ 0&0&0&a_{3}\end{bmatrix}\,.$ (40)
@@
$\begin{bmatrix}1\\ a_{1}\\ a_{2}\\ a_{3}\end{bmatrix}=\begin{bmatrix}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&1&-1\\ -1&1&1&-1\end{bmatrix}\begin{bmatrix}e_{1}\\ e_{2}\\ e_{3}\\ e_{4}\end{bmatrix}\,.$ (41)
@@
$\begin{bmatrix}e_{1}\\ e_{2}\\ e_{3}\\ e_{4}\end{bmatrix}=\frac{1}{4}\begin{bmatrix}1&1&1&-1\\ 1&1&-1&1\\ 1&-1&1&1\\ 1&-1&-1&-1\end{bmatrix}\begin{bmatrix}1\\ a_{1}\\ a_{2}\\ a_{3}\end{bmatrix}\,,$ (42)
@@
so that positivity and normalization of the sum of the eigenvalues allow to represent the set of all Bell-diagonal states parameterized by the three-dimensional vector $(a_{1},a_{2},a_{3})$ as shown in Fig. (1).
V.2 Werner states
@@
A particular subclass of the Bell-diagonal states are the Werner states $\rho_{w}$, described by one free parameter $|a_{1}|=|a_{2}|=|a_{3}|\equiv w$. Formally, a Werner state is a convex combination of the maximally mixed state with any one of the four Bell states. This class of states, first introduced by R. F. Werner Werner1989 , is important as it provides the paramount example of a set of states that for $\frac{1}{3}\leq w\leq\frac{1}{\sqrt{2}}$ are entangled but nevertheless local.
V.3 CHSH-local states and their properties
Let $(d_{1},d_{2},d_{3})$ be the singular values in the decomposition Eq. (35) and, without loss of generality let $d_{1}\geq d_{2}\geq d_{3}$. As proved in Ref. Horodecki1995 , a state is CHSH-local, i.e. it does not violate the CHSH inequality Eq. (3), if and only if
$d_{1}^{2}+d_{2}^{2}\leq 1.$ (43)
We prove the following:
Lemma 1. If $\rho$ is a CHSH-local state with given $(\alpha_{0i},\alpha_{i0},A)$ then the state $\rho^{-}$ given by $(-\alpha_{0i},-\alpha_{i0},A)$ is also CHSH-local.
Proof. It suffices to consider a local transformation that for $i=1,2,3$ changes $\sigma_{i}$ into $-\sigma_{i}$.
Lemma 2. For any CHSH-local state $\rho$ with given $(\alpha_{0i},\alpha_{i0},A)$ there exists a CHSH-local state $\rho^{\prime}$ with the same $A$ but with $\alpha_{0i}=\alpha_{i0}=0$
Proof. Consider the convex combination
$\rho^{\prime}=\frac{1}{2}\rho+\frac{1}{2}\rho^{-}.$ (44)
Indeed, for such a combination we have
$\alpha=\begin{bmatrix}\alpha_{00}&0\\ 0&A\end{bmatrix}.$ (45)
Now, convex combinations of local states are local: by virtue of the property of shared randomness, if there exists a hidden variable model for each of the states in the combination, then there must exist a hidden variable model for their classical statistical mixture.
V.4 Properties of symmetric states
Proposition 1. For any jointly convex and unitarily invariant functional $D$ of pairs of two-qubit density matrices
$\min_{\rho^{loc}\in\sigma}D(\rho_{w},\rho^{loc})=\min_{\rho^{loc}_{w}\in\sigma}D(\rho_{w},\rho^{loc}_{w})\,,$ (46)
where $\rho_{w}$ is a Werner state, $\rho^{loc}$ is a CHSH-local state, and $\rho^{loc}_{w}$ is a CHSH-local Werner state.
Proof. The proof is based on the following observations:
  • Obs. 1. Invariance of Werner states under the same local unitary transformations on both subsystems:
    $\forall\ {\rm unitary}\ U:\qquad\left(U\otimes U\right)\rho_{w}\left(U^{\dagger}\otimes U^{\dagger}\right)=\rho_{w}.$ (47)
  • Obs. 2. Local unitary transformations are resource non-generating (RNG) for nonlocality; in particular, $\left(U\otimes U\right)\rho^{loc}\left(U^{\dagger}\otimes U^{\dagger}\right)$ remains local.
  • Obs. 3. Any convex combination of local states is a local state.
  • Obs. 4. Twirling always projects on Werner states. Therefore, we have that for any given class of density matrices $\rho^{a}$
[…]
    where the integral is taken over the Haar measure on the set of unitary transformations.
From the unitary invariance of $D$ we have
$D(\rho_{w},\rho^{loc})$ $=$ $D\left(U\otimes U\rho_{w}U^{\dagger}\otimes U^{\dagger}\,,\,U\otimes U\rho^{loc}U^{\dagger}\otimes U^{\dagger}\right)$ (49)
[…]
$=$ $D(\rho_{w},\rho^{loc}_{w})\,.$ (52)
Here the first and second lines come from unitary invariance of $D$; the third follows from the joint convexity; and the fourth from Obs. 1 - 4 above.
Similar propositions hold for general Bell-diagonal states; however, in the general case the proof holds for fairly smaller symmetry classes.
Proposition 2. For any jointly convex and unitarily invariant functional $D$ of pairs of two-qubit density matrices, the closest CHSH-local state to a Bell-diagonal state is Bell-diagonal:
$\min_{\rho_{loc}\in\sigma}f(\rho_{BD},\rho_{loc})=\min_{\rho_{BD,loc}\in\sigma}f(\rho_{BD},\rho_{BD,loc})\,,$ (53)
where $\rho_{BD}$ denotes a Bell-diagonal state, $\rho_{loc}$ a CHSH-local state, and $\rho_{BD,loc}$ a CHSH-local Bell-diagonal state.
Proof. The symmetry class that we consider contains the simultaneous local $\pi$-rotations of qubits around the three Pauli axes. Rotation $R_{1}$ around $\sigma_{1}$ changes $\sigma_{2}$ and $\sigma_{3}$ to $-\sigma_{2}$ and $-\sigma_{3}$ respectively. Rotations $R_{2}$ and $R_{3}$ operate analogously. These transformations are local and unitary, therefore they cannot change the locality of a quantum state. Moreover, these transformations preserve Bell-diagonal states. Indeed, let us consider the general state representation Eq. (31). Under rotation $R_{1}$ we have the following transformation:
$\alpha=\begin{pmatrix}1&a_{01}&a_{02}&a_{03}\\ a_{10}&a_{11}&a_{12}&a_{13}\\ a_{20}&a_{21}&a_{22}&a_{23}\\ a_{30}&a_{31}&a_{32}&a_{33}\end{pmatrix}\,\,\longrightarrow\,\,\alpha_{1}=\begin{pmatrix}1&a_{01}&-a_{02}&-a_{03}\\ a_{10}&a_{11}&-a_{12}&-a_{13}\\ -a_{20}&-a_{21}&a_{22}&a_{23}\\ -a_{30}&-a_{31}&a_{32}&a_{33}\end{pmatrix}.$ (54)
$\alpha_{1}^{\prime}=\begin{pmatrix}1&a_{01}&0&0\\ a_{10}&a_{11}&0&0\\ 0&0&a_{22}&a_{23}\\ 0&0&a_{32}&a_{33}\end{pmatrix}.$ (55)
Applying rotation $R_{2}$ to $\rho^{\prime}$ and averaging the result with $\rho^{\prime}$ leads to the state with diagonal representation Eq. (39), i.e. a Bell-diagonal state. The total transformation that projects a given state to a Bell-diagonal state takes the compact form
$\rho\,\,\longrightarrow\,\,\frac{1}{2}\left(\frac{1}{2}(\rho+R_{1}(\rho))+R_{2}\left(\frac{1}{2}(\rho+R_{1}(\rho))\right)\right).$ (56)
This transformation obviously preserves the Bell-diagonal states, does not change the locality, and is a composition of joint local unitaries and averaging. The proof of Proposition 2 is thus immediately completed by straightforwardly adapting the reasoning followed in the proof of Proposition 1.
VI CHSH nonlocality of two-qubit states
Exploiting the results derived in the previous section, we can proceed to quantify the CHSH nonlocality of two-qubit states using the Hilbert space based distance and relative entropy measures with respect to the set of CHSH-local two-qubit states. Explicit examples will include Werner states, Bell-diagonal states, and reduced convex combinations of Bell states. These important classes of states are routinely considered in the study of quantum resource theories, starting from the pioneering investigations on entanglement and discord Wootters1997 ; Wootters1998 ; Henderson_Vedral2001 .
VI.1 Werner states
VI.1.1 Hilbert-Schmidt distance
Consider the squared Hilbert-Schmidt (HS) distance defined on a generic pair $\{\rho_{1},\rho_{2}\}$ of two-qubit states as induced by the HS norm Eq. (20). Using Eq. (31) and the properties of the Pauli matrices:
$\|\rho_{1}-\rho_{2}\|_{HS}\,=\,\|\alpha_{1}-\alpha_{2}\|_{HS}\,.$ (57)
It follows from Proposition 1, Eq. (46), that the closest CHSH-local state to a Werner state is also a Werner state with $w\leq 1/\sqrt{2}$. Hence, for CHSH-nonlocal Werner states, i.e. with $w\geq 1/\sqrt{2}$ :
$\mathcal{M}_{HS}(\rho_{w})=\min_{\rho^{loc}_{w}\in\sigma}\left\|\rho_{w}-\rho^{loc}\right\|_{HS}^{2}=\left\|\frac{1}{\sqrt{2}}\mathbf{1}-w\mathbf{1}\right\|_{HS}^{2}=3\left(\frac{1}{\sqrt{2}}-w\right)^{2}\,,$ (58)
where $\mathbf{1}$ is the identity matrix and $w\in\left(1/\sqrt{2},1\right]$. Correctly, the maximum is achieved by the maximally entangled Bell states corresponding to $w=1$ :
$\mathcal{M}_{HS}^{max}\,=\,\frac{3}{2}\left(3-2\sqrt{2}\right)\,.$ (59)
The behavior of the normalized HS geometric measure of CHSH nonlocality for Werner states ~$\widetilde{\mathcal{M}}_{HS}=\mathcal{M}_{HS}(\rho_{w})/\mathcal{M}_{HS}^{max}$ as a function of the $w$ parameter is reported in Fig. 2. The plot is a parabola, as the the HS distance is simply the Euclidean distance in the three-dimensional space of Fig. 1.
VI.1.2 Hellinger distance
$\mathcal{M}_{He}(\rho_{w})=\min_{\rho_{w}^{loc}\in\sigma}D^{2}_{He}(\rho_{w},\rho_{w}^{loc})=2-\frac{1}{2}\left(3\sqrt{(1-w)\left(1-\frac{1}{\sqrt{2}}\right)}+\sqrt{(1+3w)\left(1+\frac{3}{\sqrt{2}}\right)}\right)\,,$ (60)
where $w\in(1/\sqrt{2},1]$. Here we have used the relation between the eigenvalues of the state density matrix and the Werner parameter $w$ which is derived directly from Eq. (42). As $|a_{1}|=|a_{2}|+|a_{3}|=w$, the signs of the matrix elements $a_{i}$ identify the states corresponding to different corners of the state tetrahedron in Fig. 1. The Hellinger measure is maximized by the Bell states at $w=1$ :
$\mathcal{M}_{He}^{max}=2-\sqrt{1+\frac{3}{\sqrt{2}}}\,\,.$ (61)
The behavior of the normalized Hellinger geometric measure of CHSH nonlocality for Werner states ~$\widetilde{\mathcal{M}}_{He}=\mathcal{M}_{He}(\rho_{w})/\mathcal{M}_{He}^{max}$ as a function of the $w$ parameter is shown in Fig. 2.
VI.1.3 Bures distance
In the case of commuting operators the Bures distance and the Hellinger distance coincide. This is indeed the case for Werner density matrices. Therefore, the Bures and Hellinger measures of CHSH nonlocality coincide on Werner states $\rho_{w}$:
$\mathcal{M}_{Bu}(\rho_{w})=\mathcal{M}_{He}(\rho_{w})\,.$ (62)
VI.1.4 Trace distance
$\mathcal{M}_{Tr}(\rho_{w})=\min_{\rho_{w}^{loc}\in\sigma}D_{Tr}(\rho_{w},\rho_{w}^{loc})=\frac{3}{2}\left|w-\frac{1}{\sqrt{2}}\right|=\frac{3}{2}\left(w-\frac{1}{\sqrt{2}}\right)\,,$ (63)
where the last equality follows from the fact that without loss of generality we can limit the analysis to the interval $w\in(1/\sqrt{2},1]$. The maximum value is achieved at $w=1$ by the Bell states:
$\mathcal{M}_{Tr}^{max}=\frac{3}{2}\left(1-\frac{1}{\sqrt{2}}\right)\,.$ (64)
The normalized trace measure of CHSH nonlocality for Werner states ~$\widetilde{\mathcal{M}}_{Tr}=\mathcal{M}_{Tr}(\rho_{w})/\mathcal{M}_{Tr}^{max}$ as a function of the $w$ parameter is shown in figure 2.
VI.1.5 Relative entropy
It is straightforward to verify that for a nonlocal Werner state in the interval $w\geq 1/\sqrt{2}\,$ its relative entropy Eq. (14) with respect to the set $\sigma$ of local states is minimized by the local state with $w=1/\sqrt{2}$ , so that:
$\mathcal{M}_{Re}(\rho_{w})=\min_{\rho_{w}^{loc}\in\sigma}S(\rho_{w}|\rho_{w}^{loc})$ $=$ $\frac{3(1-w)}{4}\log_{2}\left(\frac{3(1-w)}{4}\right)+\frac{1+3w}{4}\log_{2}\left(\frac{1+3w}{4}\right)+$ (65)
The maximum value achieved by the Bell states at $w=1$ reads
$\mathcal{M}_{Re}^{max}=2-\log_{2}{\bigg{(}1+\frac{3}{\sqrt{2}}\bigg{)}}\,.$ (66)
The normalized relative entropy measure of CHSH nonlocality for Werner states ~$\widetilde{\mathcal{M}}_{Re}=\mathcal{M}_{Re}(\rho_{w})/\mathcal{M}_{Re}^{max}$ as a function of the $w$ parameter is reported in Fig. 2.
Figure 2: Normalized geometric measures of CHSH nonlocality for Werner states as a function of the Werner parameter $w$ in the interval $\frac{1}{\sqrt{2}}\leq w\leq 1$. Black dashed straight line: Trace measure ~$\widetilde{\mathcal{M}}_{Tr}$. Grey dashed curve: Hilbert-Schmidt measure ~$\widetilde{\mathcal{M}}_{HS}$. Black solid curve: Relative entropy measure ~$\widetilde{\mathcal{M}}_{Re}$. Grey solid curve: Hellinger measure ~$\widetilde{\mathcal{M}}_{He}$. All plotted quantities are dimensionless.
VI.2 Bell-diagonal states
VI.2.1 Hilbert-Schmidt distance
All Bell-diagonal states can be represented as points in a tetrahedron, as shown in Fig. 1, with vertices representing Bell states. The points are given by three dimensional vectors $\vec{\alpha}$ that appear in Eq. (57). This equation implies that the Euclidean distance in the space represented in Fig. 1 coincides with the HS distance in the space of density matrices of Bell-diagonal states. Moreover, for all Bell-diagonal states the closest local state is a Bell-diagonal state, as proved in Proposition 2, Eq. (53). The set of local states is represented by points in the solid figure inscribed inside the tetrahedron, see Fig. 1. Hence, the Euclidean distance between the points representing nonlocal Bell-diagonal states and the boundary of the solid figure yields the HS-based geometric measure of CHSH nonlocality. Exploiting the Horodecki locality criterion Eq. (43), the separation surface between local and nonlocal Bell-diagonal state is identified by the solutions of the equation
$\max_{i,ji\neq j}\{a_{i}^{2}+a_{j}^{2}\}=1\,.$ (67)
@@ VI.2.1 Hilbert-Schmidt distance
$\min_{\rho_{BD}^{loc}\in\sigma}\left\|\rho_{BD}-\rho_{BD}^{loc}\right\|_{HS}^{2}=\min_{a_{i}^{loc}}\frac{1}{2}\displaystyle\sum_{i=1}^{3}(a_{i}-a_{i}^{loc})^{2}\,,$ (68)
with the provision that the states $\rho_{BD}$ belong to the surface Eq. (67).
Considering first the case $|a_{1}^{loc}|,|a_{2}^{loc}|>|a_{3}^{loc}|$, we can proceed to determine the minimum by introducing a Lagrange multiplier $\lambda$. The associated Lagrange variational function reads
@@
Vanishing of the gradient of $\Gamma$ leads to
$\begin{cases}(a_{1}-a_{1}^{loc})+\lambda a_{1}^{loc}=0,\\ (a_{2}-a_{2}^{loc})+\lambda a_{2}^{loc}=0,\\ (a_{3}-a_{3}^{loc})=0,\\ (a_{1}^{loc})^{2}+(a_{2}^{loc})^{2}-1=0.\end{cases}$
@@
$a_{1}^{loc}=\pm\frac{a_{1}}{\sqrt{a_{1}^{2}+a_{2}^{2}}}{2},$
$a_{2}^{loc}=\pm\frac{a_{2}}{\sqrt{a_{2}^{2}+a_{3}^{2}}},$
@@
$\mathcal{M}_{HS}(\rho_{BD})=\frac{3}{4}\sum\limits_{i=x,y,z}(a_{i}-a_{i}^{loc})^{2}=\frac{3}{4}\left(\sqrt{a_{1}^{2}+a_{2}^{2}}\mp 1\right)^{2}$ (71)
whose minimum clearly correspond to the minus sign (the plus sign in (71)), since $t_{x}^{2}+t_{y}^{2}\geq 1$. The remaining cases $|a_{2}^{loc}|,|a_{3}^{loc}|>|a_{1}^{loc}|$ and $|a_{1}^{loc}|,|a_{3}^{loc}|>|a_{2}^{loc}|$ can be immediately found by symmetry and the global minimum is determined over all possible cases. For the maximally nonlocal Bell states the minimum is attained when $|a_{1}^{loc}|=|a_{2}^{loc}|=|a_{3}^{loc}|\equiv w^{2}$, with $w$ the Werner parameter. The constraints in Eq. (69) thus reduce to $12w^{2}=$ 1, i.e. $w=1/\sqrt{2}$, and $\mathcal{M}_{HS}$ achieves its maximum $\mathcal{M}_{HS}^{max}$, Eq. (59). Note that this analytical expression is not always valid, as no constraint has been imposed on the positivity of the density operator associated with the set $\{a_{i}^{loc}\}$. Therefore, in order to use expression (71), one must also take such constraints into account. If the local state does not satisfy these constraints, the distance must be computed numerically. Using Eqs. (41) and (42) the transformation from the matrix elements $\{a_{i}\}$ to the eigenvalues $\{e_{i}\}$ of the density matrix follows immediately. The behavior of the normalized HS measure of CHSH nonlocality ~$\widetilde{\mathcal{M}}_{HS}=\mathcal{M}_{HS}/\mathcal{M}_{HS}^{max}$ for Bell-diagonal states as a function of the eigenvalues $e_{1}$, $e_{2}$ in the case of $e_{4}=0$ is reported in Fig. 3.
Figure 3: Normalized Hilbert-Schmidt measure ~$\widetilde{\mathcal{M}}_{HS}$ of CHSH nonlocality for Bell-diagonal states as a function of the two free parameters $e_{1}$ and $e_{2}$, with $e_{4}=0$, $e_{3}=1-e_{1}-e_{2}$, and normalization coefficient $\mathcal{M}_{HS}^{max}$. The maximum ~$\widetilde{\mathcal{M}}_{HS}=1$ is reached in the three cases $(e_{1},e_{2},e_{3})=(1,0,0)$; $(e_{1},e_{2},e_{3})=(0,1,0)$; and $(e_{1},e_{2},e_{3})=(0,0,1)$. Each case corresponds to a different Bell state. The central region of minima is composed by local states, i.e. states with ~$\widetilde{\mathcal{M}}_{HS}=0$. All plotted quantities are dimensionless.
VI.2.2 Hellinger distance
Each of the eigenvalues $\{e_{i}^{k}\}$ of a density matrix $\rho_{k}$ is the probability that a system in state $\rho_{k}$ collapses in the eigenstate relative to that eigenvalue. For a Bell-diagonal state $\rho_{BD}$ the eigenstates are obviously the four Bell states. Recalling Proposition 2, Eq. (53), and Eq. (23) the minimization process on the Hellinger distance squared yields the Hellinger geometric measure of CHSH nonlocality $\mathcal{M}_{He}(\rho_{BD})$ for Bell-diagonal states:
$\mathcal{M}_{He}(\rho_{BD})=\min_{\rho_{BD}^{loc}\in\sigma}D^{2}_{He}(\rho_{BD},\rho_{BD}^{loc})=\min_{\rho^{loc}\in\sigma}\left\{\displaystyle\sum_{i=1}^{4}\bigg{(}\sqrt{e_{i}}-\sqrt{e_{i}^{loc}}\bigg{)}^{2}\right\}\,.$ (72)

### v5

@@ A.1 Bell-diagonal states
$\rho_{\mathrm{BD}}=\frac{1}{4}\bigg[\mathds{1}_{2}\otimes\mathds{1}_{2}+\displaystyle\sum_{i=1}^{3}a_{i}\sigma_{i}\otimes\sigma_{i}\bigg]\,$, (67)
where $\mathds{1}_{2}$ denotes the identity matrix in two dimensions. Matrix $\alpha$ of Eq. (27) can then be recast in the form
$\alpha_{\mathrm{BD}}=\left[\begin{array}[]{cccc}1&0&0&0\\ 0&a_{1}&0&0\\ 0&0&a_{2}&0\\ 0&0&0&a_{3}\end{array}\right]\,.$ (68)
@@
$\left[\begin{array}[]{c}1\\ a_{1}\\ a_{2}\\ a_{3}\end{array}\right]=\left[\begin{array}[]{cccc}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&1&-1\\ -1&1&1&-1\end{array}\right]\left[\begin{array}[]{c}e_{1}\\ e_{2}\\ e_{3}\\ e_{4}\end{array}\right]\,.$ (69)
@@
$\left[\begin{array}[]{c}e_{1}\\ e_{2}\\ e_{3}\\ e_{4}\end{array}\right]=\frac{1}{4}\left[\begin{array}[]{cccc}1&1&1&-1\\ 1&1&-1&1\\ 1&-1&1&1\\ 1&-1&-1&-1\end{array}\right]\left[\begin{array}[]{c}1\\ a_{1}\\ a_{2}\\ a_{3}\end{array}\right]\,.$ (70)
@@
so that the positivity and normalization of the sum of the eigenvalues allow us to represent the set of all Bell-diagonal states parameterized by the three-dimensional vector $(a_{1},a_{2},a_{3})$ as shown in Fig. (4).
A.2 Werner states
@@
A particular subclass of the Bell-diagonal states are the Werner states $\rho_{w}$, described by one free parameter $|a_{1}|=|a_{2}|=|a_{3}|\equiv w$. Formally, a Werner state is a convex combination of the maximally mixed state with any one of the four Bell states. This class of states, first introduced by R. F. Werner [10], is important as it provides the paramount example of a set of states that for $\frac{1}{3}\leq w\leq\frac{1}{\sqrt{2}}$ are entangled but nevertheless local.
Appendix B Bell-diangonal states
B.1 Hilbert-Schmidt distance
@@
$\min_{\rho_{\mathrm{BD}}^{loc}\in\mathcal{L}}\left\|\rho_{\mathrm{BD}}-\rho_{\mathrm{BD}}^{loc}\right\|_{2}=\min_{a_{i}^{loc}}\frac{1}{2}\sqrt{\displaystyle\sum_{i=1}^{3}(a_{i}-a_{i}^{loc})^{2}}\,$, (71)
with the provision that the states $\rho_{\mathrm{BD}}$ belong to the surface Eq. (50).
Considering first the case $|a_{1}^{loc}|,|a_{2}^{loc}|>|a_{3}^{loc}|$, we can proceed to determine the minimum by introducing a Lagrange multiplier $\lambda$. The associated Lagrange variational (squared) function reads
@@
The vanishing of the gradient of $\Gamma$ leads to
$\left\{\begin{array}[]{rcl}(a_{1}-a_{1}^{loc})+\lambda a_{1}^{loc}&=&0,\\ (a_{2}-a_{2}^{loc})+\lambda a_{2}^{loc}&=&0,\\ (a_{3}-a_{3}^{loc})&=&0,\\ (a_{1}^{loc})^{2}+(a_{2}^{loc})^{2}-1&=&0.\end{array}\right$. (73)
@@
$a_{1}^{loc}=\pm\frac{a_{1}}{\sqrt{a_{1}^{2}+a_{2}^{2}}},$
$a_{2}^{loc}=\pm\frac{a_{2}}{\sqrt{a_{2}^{1}+a_{2}^{2}}},$
@@
$\mathcal{M}_{\mathrm{HS}}(\rho_{\mathrm{BD}})=\frac{1}{2}\sqrt{\sum\limits_{i=x,y,z}(a_{i}-a_{i}^{loc})^{2}}=\frac{1}{2}\left(\sqrt{a_{1}^{2}+a_{2}^{2}}\mp 1\right)$ (75)
whose minimum clearly corresponds to the minus sign (the plus sign in (74)), since $a_{x}^{2}+a_{y}^{2}\geq 1$. It should be noted that the solution (74) is acceptable only if it satisfies the condition $|a_{1}^{loc}|,|a_{2}^{loc}|>|a_{3}^{loc}|$. The remaining cases $|a_{2}^{loc}|,|a_{3}^{loc}|>|a_{1}^{loc}|$ and $|a_{1}^{loc}|,|a_{3}^{loc}|>|a_{2}^{loc}|$ can be immediately found by symmetry, and the global minimum is determined over all possible cases, provided that the corresponding $|a_{i}^{loc}|$ are correctly ordered.
It may happen that none of the three cases gives rise to a coherent solution; in this case, it is not possible to express the HS measure in the form (75). This is the case of Werner states, where $|a_{i}|=w$, but the solution (74) implies $|a_{1}^{loc}|,|a_{2}^{loc}|=1/\sqrt{2}<|a_{3}^{loc}|$. In this case the correct solution is indeed different from (75) (see (41)).
Using Eqs. (69) and (70) the transformation from the matrix elements $\{a_{i}\}$ to the eigenvalues $\{e_{i}\}$ of the density matrix follows immediately. The behavior of the normalized HS measure of CHSH nonlocality $\widetilde{\mathcal{M}}_{\mathrm{HS}}=\mathcal{M}_{\mathrm{HS}}/\mathcal{M}_{\mathrm{HS}}^{max}$ for Bell-diagonal states as a function of the eigenvalues $e_{1}$, $e_{2}$ in the case of $e_{4}=0$ is reported in Fig. 3.
B.2 Hellinger distance
Each of the eigenvalues $\{e_{i}^{k}\}$ of a density matrix $\rho_{k}$ is the probability that a system in state $\rho_{k}$ collapses in the eigenstate relative to that eigenvalue. For a Bell-diagonal state $\rho_{\mathrm{BD}}$ the eigenstates are obviously the four Bell states. Recalling Proposition 2, Eq. (32), and Eq. (11) the minimization process on the Hellinger distance squared yields the Hellinger geometric measure of CHSH nonlocality $\mathcal{M}_{\mathrm{He}}(\rho_{\mathrm{BD}})$ for Bell-diagonal states:
$\mathcal{M}_{\mathrm{He}}(\rho_{\mathrm{BD}})=\min_{\rho_{\mathrm{BD}}^{loc}\in\mathcal{L}}D^{2}_{\mathrm{He}}(\rho_{\mathrm{BD}},\rho_{\mathrm{BD}}^{loc})=\min_{\rho^{loc}\in\mathcal{L}}\left\{\displaystyle\sum_{i=1}^{4}\bigg(\sqrt{e_{i}}-\sqrt{e_{i}^{loc}}\bigg)^{2}\right\}\,.$ (76)